What volume of 1.0 M hydrochloric acid will be required to produce 800 cm3 hydrogen gas by reacting with excess zinc? (a single replacement reaction)
1
Expert's answer
2020-08-29T14:02:03-0400

"1 \ cm^3 = 1 \ ml"

Moles of "H_2" gas in "800 \ ml" "= \frac{800}{22400}=0.03571" "(" As volume of "1" mole of gas is "22400\ ml)"

Moles of "H" atoms will be twice to that of "H_2" .

So, moles of hydrogen"=2\times 0.03571=0.07143"


Applying mole balance over "H" to reactant"( \ HCl \ )" and product side "( H_2)" ,

Molarity of "HCl" "( M)= 1 M"

So let volume required be "V" .

"\implies MV=0.07143"

"\implies 1.0\times V= 0.07143"

"\implies V = 0.07143 \ l = 71.43 \ ml= 71.43 \ cm^3"